If I=<24,36,42> be an ideal of a ring Z then find a such that I=<a>.
1
Expert's answer
2020-10-21T17:15:47-0400

by properties of the Ideal:

"\forall a,b\in I\; a-b\in I"

Substitute all possible values and find their GCD:

"36-24=12\\
42-24=18\\
42-36=6\\
GCD=6"

From the fact that 24, 36, 42 are divided by 6, it follows "I=<6>"